Rewrite the equation 2x – 9y = –72 as a function of x
Please HELP!

Answer by dolly(163)

Given 2x – 9y = –72
==> - 9y = -72 - 2x [ subtracting 2x from both the sides]
==> 9y = 72 + 2x [ multiplying by -1 on both the sides]
==> y = (72+2x)/9
Thus y = (2/9)x + 8
Best of luck!!!